OpenStudy (anonymous):

Simplify the expression: (5z^2)(4z^3)?

OpenStudy (anonymous):

a^2(bc)b^2 use this

OpenStudy (anonymous):

\[\LARGE x^a\cdot x^b=x^{a+b}\] since we have ONLY multiply ! we don't need bracket ! ...take them off \[\huge 5z^2\cdot 4z^3=20z^?\]

OpenStudy (anonymous):

so the answer would be 20z^5 then?

OpenStudy (anonymous):

thanks!!